OpenStudy (anonymous):

A drag racer starts her car from rest and accelerates at 7.7 m/s^2 for the entire distance of 518 m. a) How long did it take the car to travel this distance? Answer in units of s. b) What is its speed at the end of the run? Answer in units of m/s.

OpenStudy (anonymous):

@lostmermaid do you know the equation of uniform motion.

OpenStudy (anonymous):

Here given, u = 0 ( as the car starts from the rest) a = 7.7 m/s^2 s = 518 m (a) t = ? Use, \[s = ut +\frac{ 1 }{ 2 } a t ^{2}\]

OpenStudy (anonymous):

(b) v = u + at